OpenStudy (anonymous):

y over negative 4 + 5 = 15

OpenStudy (anonymous):

y over-4

OpenStudy (goldphenoix):

\[\large \large \frac{ y }{ -4+5 } = 15\] Does it look like that?

OpenStudy (anonymous):

y -+5=15 -4

OpenStudy (goldphenoix):

\[\large \large \frac{ y }{ -4 } + 5 = 15\] Like that?

OpenStudy (anonymous):

yeah

OpenStudy (goldphenoix):

Alright. What do you think you do first to find y?

OpenStudy (anonymous):

find y

OpenStudy (goldphenoix):

Yes. So what do you think we do first?

OpenStudy (anonymous):

+5 both side

OpenStudy (goldphenoix):

We want to add 5 to both side?

OpenStudy (anonymous):

i mean -4 both side

OpenStudy (goldphenoix):

-4 to both side? Are you sure?

OpenStudy (anonymous):

yeah

OpenStudy (goldphenoix):

Hmm. Alright... What does that give us?

OpenStudy (anonymous):

i think. what yu think

OpenStudy (goldphenoix):

Well. If you want to find any variable, you want to do inverese order of operation. Yes, we do subtract. But why 4? Tha doesn't lead us to anywhere. If we subtract it by 5, then that would lead us to somewhere.

OpenStudy (goldphenoix):

So if you subtract 5 by both side, then 5-5 = ? 15 - 5 =?

OpenStudy (anonymous):

what i got to do with does number 10 in 0

OpenStudy (goldphenoix):

What do you mean?

OpenStudy (anonymous):

5-5=0 15-5=10 what else i need to do

OpenStudy (goldphenoix):

Oh wait. So now you have: \[\large \large \frac{ y }{ -4 } = 10\] We want to find y, you know how?

OpenStudy (anonymous):

so 5 in 15 gone

OpenStudy (goldphenoix):

Yes. Since, 15 - 5 = 10. Do you know what to do next?

OpenStudy (anonymous):

15-5=10 -5 -5=10 10=10

OpenStudy (anonymous):

is tha right

OpenStudy (goldphenoix):

Nope. Here, another way of writing that fraction is y / -4. Correct?

OpenStudy (anonymous):

yeah

OpenStudy (goldphenoix):

So we want to do inverse operation to cancel out -4. So what's the opposite of division?

OpenStudy (anonymous):

im confuse

OpenStudy (goldphenoix):

What are you confused with?

OpenStudy (anonymous):

how we suppose to cancel out -4 in both side

OpenStudy (goldphenoix):

Like I said. You want to do inverse operation. So what's the opposite of division?

OpenStudy (anonymous):

subtraction

OpenStudy (goldphenoix):

o.O. Hmm. It's not.

OpenStudy (goldphenoix):

Let say you have 10 - x = 4. What would you do to find x?

OpenStudy (anonymous):

*

hero (hero):

Addition and subtraction are opposites of each other. Multiplication and division are reciprocals of each other. Both opposites and reciprocal operations are considered in the class of inverse operations.

hero (hero):

There's no such thing as "opposite of division"

OpenStudy (goldphenoix):

Correct. You should remember tat, doulo. You will need that later on. So you want to do multiplication, correct, doulo?

OpenStudy (goldphenoix):

Thanks, Hero. I will keep that in mind.

OpenStudy (anonymous):

6=x

OpenStudy (goldphenoix):

Nope. Don't go ahead of me. We want to do multiplication, yes or no?

OpenStudy (anonymous):

no we subtraction

OpenStudy (goldphenoix):

Do you remembered what Hero said?

OpenStudy (anonymous):

so im suppose to do division 10 in 4

OpenStudy (goldphenoix):

No-no. y / -4 is another way of writing it. To get rid of -4, you want to do inverse operation. So what's the 'reciprocal' of division?

OpenStudy (anonymous):

y count as 1 right

OpenStudy (goldphenoix):

What do you mean? y can be anything.

OpenStudy (anonymous):

so can we get rid of -4

OpenStudy (goldphenoix):

Yes!!! :| Like I said, we want to do inverse operation to get rid of -4. So what's the 'reciprocal' of division?

OpenStudy (anonymous):

y=4+5=15

OpenStudy (goldphenoix):

Please answer my question. What's the 'reciprocal' of division? You're only 2 steps away from answering your question!!!

OpenStudy (anonymous):

multiplication right?

OpenStudy (marissalovescats):

Hi @doulo I hope this helps you understand form out messages